OpenStudy (anonymous):

You have a rectangular pot measuring 16feet by 25feet in a community garden. You want to grow tomato plants that each need 8 square feet of space and pepper plants that each need 5 square feet. Write an equation that models how many pepper plants you can grow. How many pepper plants can you grow if you grow 15 tomato plants? All in standard form....Has it stumped anyone else??

OpenStudy (mertsj):

5p=16(25)

OpenStudy (mertsj):

8(15)+5p=16(25)

OpenStudy (anonymous):

oh my you're brilliant!